# Acknowledgments and Disclosure of Funding

### v2

@@ Abstract
Large language models (LLMs) can be improved by aligning with human preferences through fine-tuning—the so-called reinforcement learning from human feedback (RLHF). However, the cost of fine-tuning an LLM is prohibitive for many users. Due to their ability to bypass LLM fine-tuning, prediction-time tokenwise reward-guided text generation (RGTG) methods have recently been proposed. They use a reward model trained on full sequences to score partial sequences during decoding in a bid to steer the generation towards sequences with high rewards. However, these methods have so far been only heuristically motivated and poorly analyzed. In this work, we show that reward models trained on full sequences are not compatible with scoring partial sequences. To alleviate this issue, we propose to train a Bradley-Terry reward model on partial sequences explicitly, and autoregressively sample from the implied tokenwise policy during decoding time. We study the properties of this reward model and the resulting policy: we show that this policy is proportional to the ratio of two distinct RLHF policies. Our simple approach outperforms previous RGTG methods and performs similarly to strong offline baselines without large-scale LLM finetuning.
@@ 1 Introduction
Large language models (LLMs) provide a modern foundation for most, if not all, text generation tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b). In practice, significant improvements in the quality of text generation are achieved by aligning LLMs to human preferences (Stiennon et al., 2020b; Ouyang et al., 2022). This is typically performed via reinforcement learning from human feedback (RLHF), which involves two steps: i) learning a reward model from preference data and ii) fine-tuning an LLM to maximize expected rewards by reinforcement learning (Ziegler et al., 2019b). Usually, this is done via a reinforcement learning algorithm such as proximal policy optimization (PPO, Schulman et al., 2017). Nevertheless, recently, Rafailov et al. (2023) showed that the reward modeling step (i) can be bypassed by directly fine-tuning an LLM with preference data, resulting in a method called direct preference optimization (DPO). While this simplifies RLHF, the fine-tuning step (ii) remains prohibitively costly for most users since it requires high-performance computational resources with large GPUs.
@@
In order to alleviate the computational issue above, Khanov et al. (2024); Deng & Raffel (2023) explored tokenwise reward-guided text generation (RGTG) techniques that avoid any fine-tuning of the LLM. More precisely, the LLM remains frozen (i.e., not finetuned), and the reward model is used at the decoding time to adjust the softmax scores of the tokens directly. Unlike DPO, this line of work retains the reward modeling step, but training reward models are typically a much cheaper endeavor than finetuning text-generation LLMs since smaller models can be utilized for reward modeling. Furthermore, reward models are modular: they can easily be composed and reused without any cost to guide text generation in conjunction with any base LLM. In contrast, RLHF via DPO requires fine-tuning of every LLM that we wish to improve based on human preference data.
@@
While RGTG is an interesting alternative to the standard offline RLHF, it is often based on heuristics and still poorly analyzed. For instance, ARGS (Khanov et al., 2024) proposed to simply use a reward model trained on full sequences to score each partial sequence during autoregressive decoding. Meanwhile, Deng & Raffel (2023) used a custom tokenwise loss to distill a reward model trained on full sequences. Thus, it is unclear if these approaches can give rise to a sound tokenwise text generation policy. Controlled decoding (CD; Mudgal et al., 2024), on the other hand, uses rollouts from the base model along with a reward model trained on full sequences to distill the partial reward.
@@
In this work, we analyze this common RGTG approach. First, we show that the usage of full-sequence reward models for scoring partial sequences in a tokenwise policy is pathological. To alleviate this, we propose to explicitly train a Bradley-Terry reward model on partial sequences. We prove that this text generation policy is a ratio of two different RLHF policies trained on sequences of different lengths. Ideally, the policy would be derived from a single RLHF policy, but as we shall also show in Section 3, the exact computation of such a policy is intractable. By deriving the policy from a ratio of distinct RLHF policies, we obtain a tractable sampling procedure. This is akin to the argument of Zhao et al. (2024) where they use a similar ratio but for deriving a sequential Monte Carlo method.
@@
    We thus propose to explicitly train a Bradley-Terry reward model on partial sequences and sample from the induced per-token policy induced by it during the decoding time.
@@
    We show that this reward model induces a ratio of two distinct RLHF policies over sequences with different lengths. This is a tradeoff that one must make to make tokenwise RGTG free of the pathology and yet still tractable.
@@
    Extensive experiments with recent LLMs and various text generation tasks validate our insights.
@@ 2.1 Reinforcement Learning from Human Feedback
Given a preference dataset $\mathcal{D}=\{({\mathbf{x}}_{k},{\mathbf{y}}_{wk},{\mathbf{y}}_{lk})\}_{k=1}^{K}$ containing $K$ triples of token sequences $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$, Ziegler et al. (2019b) and Ouyang et al. (2022) proposed a technique based on reinforcement learning (RL) to align an LLM with the preference dataset. They train a parametric reward model $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$ that assigns a higher score to the “winning” (i.e., preferred) utterance ${\mathbf{y}}_{w}$ and a lower score to the “losing” utterance ${\mathbf{y}}_{l}$. This is done via the Bradley-Terry model (Bradley & Terry, 1952) which minimizes the loss:
@@
where $\sigma$ is the logistic function. Note that $r_{\phi}$ is trained to score entire utterances ${\mathbf{y}}$. Once $r_{\phi}$ is trained, it can be used to infer the probability of generating sequence ${\mathbf{y}}$ in response to ${\mathbf{x}}$, i.e., $P_{\phi}({\mathbf{y}}|{\mathbf{x}})=\nicefrac{{\exp(r_{\phi}({\mathbf{y}}|{\mathbf{x}}))}}{{\sum_{{\mathbf{y}}^{\prime}}\exp(r_{\phi}({\mathbf{y}}^{\prime}|{\mathbf{x}})}}$. Given a reference LLM, we denote by $\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})$ the conditional probability that it will generate response ${\mathbf{y}}$ to prompt ${\mathbf{x}}$ (also referred to as policy). We refer to the LLM and its policy interchangeably. One can then copy the LLM and finetune it to maximize
@@
where the KL term forms a regularizer that ensures that the finetuned model will not differ too much from the reference model. The above optimization problem can be optimized by many RL techniques, including the popular proximal policy optimization (PPO) algorithm (Schulman et al., 2017). This RL optimization is quite costly in practice due to the size of the LLM.
The above optimization (2) has a closed form solution of the form (Peters & Schaal, 2007)
@@
Maximizing the above objective with respect to $\theta$ directly finetunes the LLM without the need to learn a reward model. Furthermore, this maximization is done by supervised learning, which is generally simpler than RL. This approach, known as direct preference optimization (DPO, Rafailov et al., 2023), reduced the cost of RLHF while ensuring that the same finetuned LLM is obtained as RLHF based on PPO. Note, however, that both PPO and DPO-based RLHF are still very costly in practice since they require finetuning (a copy of) the target LLM $\pi_{\text{ref}}$.
@@ 2.2 Reward-Guided Text Generation
Note however that it is unclear whether the resulting distribution is equivalent/approximating the RLHF policy in (3). Khanov et al. (2024) do train the reward model with the Bradley-Terry loss, but it is trained only with complete sequences, i.e. $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$, while it is used to score partial sequences, i.e. $r_{\phi}({\mathbf{y}}^{1:i}|{\mathbf{x}})$. Hence, it is unclear whether the inferred scores for partial sequences are reasonable. In Section 3 we show that reward models trained only with complete sequences can assign arbitrary scores to partial sequences and in Section 5 we show empirically that the resulting RGTG policy therefore underperforms that of PPO or DPO. Meanwhile, Deng & Raffel (2023) learn the reward model by minimizing a cumulative squared loss to distill a full-sequence reward model instead of using the Bradley-Terry loss (1), making the connection to RLHF policy looser. Mudgal et al. (2024) also distill a partial reward model from a full-sequence reward model, but the tokenwise policy is not the marginal of the full-sequence policy. Nevertheless, tokens are sampled from a different tokenwise RL formulation that follows a similar derivation as RLHF.
@@
The generated sequences ${\mathbf{y}}$ are then approximately equal to the sequences generated by the RLHF policy (3). However, their method is general and does not specifically target RGTG—indeed, Zhao et al. (2024) focused on using the implied approximation of the partition function $Z({\mathbf{x}})$.
Finally, Rafailov et al. (2024) modifies DPO to obtain a partial-sequence reward model
@@
Similar to the sequence-based DPO, this reward model is then used to obtain a per-token loss function to finetune the LLM and thus, while defining a partial-sequence reward model, is not a RGTG method.
@@ 3 Pitfalls of RGTG and How to Fix Them
First, we start by analyzing the partial sequence rewards inferred from a reward model trained with full sequences only. Proof in Appendix A.
@@
A reward model $r$ trained to minimize the Bradley-Terry loss (1) on full sequences ${\mathbf{y}}^{1:|{\mathbf{y}}|}$ may assign arbitrary rewards to partial sequences ${\mathbf{y}}^{1:i}$ (where $i<|{\mathbf{y}}|$). More precisely, $r({\mathbf{y}}^{1:i}|{\mathbf{x}})=v_{{\mathbf{x}},{\mathbf{y}}^{1:i}}$ where $v_{{\mathbf{x}},{\mathbf{y}}^{1:i}}\in\mathbb{R}$ can be any value.
@@
This leads to an unidentifiability problem—see Fig. 1 for an example. If we learn a reward model based on preferences over full sequences only as proposed by Khanov et al. (2024) and Deng & Raffel (2023), then we may not obtain adequate rewards for partial sequences. As a concrete example, suppose $r$ is a reward model such that (Fig. 1)
@@
This reward model satisfies the identity in (8) and therefore could be the solution when minimizing the Bradley-Terry loss (1). If we use this reward model to sample from the induced RLHF optimal policy in (3), then the token level sampling distribution is the same as for the reference LLM $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ for all tokens except the last one. This is problematic since RLHF generally changes the token level distribution at each position, not just the last token. Hence the ARGS method (Khanov et al., 2024) may utilize a reward model that does not score partial sequences properly, which in turn may negatively impact token-by-token generation.
Figure 2: The proposed approach to alleviating the problem in Theorem 1. First, we (i) train the reward model $r$ on partial sequences explicitly (when $|{\mathbf{y}}_{w}|\neq|{\mathbf{y}}_{l}|$, we pad to the longest sequence) and (ii) sample from the weighted sum of the logits and the rewards of the next token during decoding. This is in contrast to previous RGTG methods where the reward model is trained on full sequences, but the decoding relies on partial sequence scoring (Khanov et al., 2024; Deng & Raffel, 2023).
To alleviate this issue, we propose to explicitly train the reward model with partial sequences—still using the Bradley-Terry model in contrast to Deng & Raffel (2023)—as follows (Fig. 2). We create a separate loss function for all prefix lengths $i$:
@@
Then, given that full sequence ${\mathbf{y}}_{w}$ is preferred to full sequence ${\mathbf{y}}_{l}$, we assume that the partial sequence ${\mathbf{y}}_{w}^{1:i}$ is also preferred to the partial sequence ${\mathbf{y}}_{l}^{1:i}$. Strictly speaking, it is hard for human annotators to compare partial sequences due to their incomplete nature and most preference datasets do not include preferences over partial sequences. Nevertheless, we can interpret ${\mathbf{y}}_{w}^{1:i}$ as the prefix of a winning sequence that is preferred over a losing sequence with prefix ${\mathbf{y}}_{l}^{1:i}$. The following lemma shows that the resulting reward model ensures that the probability that a first partial sequence is preferred to a second partial sequence corresponds to the probability that the first sequence is extended to a winning full sequence while the second sequence is extended to a losing full sequence according to the preference data distribution $P_{\text{data}}$. Proof in Appendix A.
@@
Hence, optimizing the partial-sequence objective (4) for all lengths $i$ determines a reward model for all response prefixes that is adequate in the sense that it induces a distribution over partial sequences that approximates the true underlying preference distribution (due to finite data) instead of assigning arbitrary rewards in the sense of Theorem 1.
@@
Algorithm 1 summarizes the decoding procedure. Contrary to the previous approach of Khanov et al. (2024), it directly follows the policy induced by the explicitly trained reward model over partial sequences. Meanwhile, compared to Deng & Raffel (2023) it leverages the standard Bradley-Terry model instead of a custom squared loss function that distills a full-sequence reward model.
Let us now analyze the tokenwise sampling distribution in (6). By the definition of conditional distributions, we can rewrite it as a ratio of two partial sequence distributions: $\pi(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})=\nicefrac{{\pi({\mathbf{y}}^{1:i}|{\mathbf{x}})}}{{\pi({\mathbf{y}}^{1:i-1}|{\mathbf{x}})}}$. It is still unclear, however, how does this distribution relates to RLHF policies—the main point of the tokenwise RGTG methods. The following theorem shows how the decoding process by following this distribution relates to RLHF-induced policies. Proof in Appendix A.
@@
Given a reward model trained according to the partial-sequence Bradley-Terry objective in (4), the induced token generation distribution $\pi$ (6) is proportional to the ratio:
@@
where $\pi_{\textnormal{RLHF},i}$ and $\pi_{\textnormal{RLHF},i-1}$ are two distinct policies over prefix sequences of length $i$ and $i-1$, respectively, induced by RLHF optimization (2).
Ideally, we would like a decoding procedure that samples the next token from a distribution that is mathematically equivalent to the conditional distribution resulting from an RLHF over full sequences. However, as shown in Theorem 3, a partial-sequence reward model $r_{\phi}$ leads to multiple RLHF decoding policies with different conditional distributions for each prefix length $i$. Hence it is not possible to have equivalence with a single RLHF policy, e.g. as obtained via PPO or DPO.
One may then ask: Which RLHF policy is best? We argue that none of them is necessarily better than the others since they simply arise from considering different prefix lengths. Note that the reward model $r_{\phi}$ leads to a distribution that approximates the true underlying preference distribution partial sequences. The problem is inherent to RLHF which takes a reference LLM with a consistent distribution over response prefixes induced by a reward model and yields different decoding policies for different prefix lengths.
@@
However, the summations in the above equation are exponentially large in the length $|{\mathbf{y}}|$ of the sequences. This exponential complexity was also noted by Zhao et al. (2024) who proposed a twisted sequential Monte Carlo technique to approximate computation and mitigate the exponential complexity. In contrast, our approach embraces the multitude of RLHF policies and leverages them in a linear time decoding procedure without any approximation of the partial sequence RLHF policies. The ratio policy (6) described here can thus be seen as a necessary tradeoff if one wants to perform tokenwise RGTG without the pathology in Theorem 1.

### v3

# A Critical Look At Tokenwise Reward-Guided Text Generation

@@ Abstract
Large language models (LLMs) can be improved by aligning with human preferences through fine-tuning—the so-called reinforcement learning from human feedback (RLHF). However, the cost of fine-tuning an LLM is prohibitive for many users. Due to their ability to bypass LLM fine-tuning, prediction-time tokenwise reward-guided text generation (RGTG) methods have recently been proposed. They use a reward model trained on full sequences to score partial sequences during decoding in a bid to steer the generation towards sequences with high rewards. However, these methods have so far been only heuristically motivated and poorly analyzed. In this work, we show that reward models trained on full sequences are not compatible with scoring partial sequences. To alleviate this, we propose to train a Bradley-Terry reward model on partial sequences explicitly, and autoregressively sample from the implied tokenwise policy during decoding. We study the properties of this reward model and the resulting policy: we show that this policy is proportional to the ratio of two distinct RLHF policies. Our simple approach outperforms previous RGTG methods and performs similarly to strong offline baselines without large-scale LLM fine-tuning. Code for our work is available at https://github.com/ahmadrash/PARGS
@@ 1 Introduction
Large language models (LLMs) provide a modern foundation for most, if not all, text generation tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b). In practice, significant improvements in the quality of text generation are achieved by aligning LLMs to human preferences (Stiennon et al., 2020b; Ouyang et al., 2022). This is typically performed via reinforcement learning from human feedback (RLHF), which involves two steps: i) learning a reward model from preference data and ii) fine-tuning an LLM to maximize expected rewards by reinforcement learning (Ziegler et al., 2019b). Usually, this is done via a reinforcement learning algorithm such as proximal policy optimization (PPO, Schulman et al., 2017). Nevertheless, recently, Rafailov et al. (2023) showed that the reward modeling step (i) can be bypassed by directly fine-tuning an LLM with preference data, resulting in a method called direct preference optimization (DPO). Although this simplifies RLHF, the fine-tuning step (ii) remains prohibitively costly for most users, since it requires high-performance computational resources with large GPUs.
@@
In order to alleviate the computational issue above, Khanov et al. (2024); Deng & Raffel (2023) explored tokenwise reward-guided text generation (RGTG) techniques that avoid any fine-tuning of the LLM. More precisely, the LLM remains frozen (i.e., not fine-tuned), and the reward model is used at the decoding time to adjust the softmax scores of the tokens directly. Unlike DPO, this line of work retains the reward modeling step, but training reward models is typically a much cheaper endeavor than fine-tuning text-generation LLMs since smaller models can be utilized for reward modeling. Furthermore, reward models are modular: they can easily be composed and reused without any cost to guide text generation in conjunction with any base LLM. In contrast, RLHF via DPO requires fine-tuning of every LLM that we wish to improve based on human preference data.
@@
While RGTG is an interesting alternative to the standard offline RLHF, it is often based on heuristics and still poorly analyzed. For instance, ARGS (Khanov et al., 2024) proposed to simply use a reward model trained on full sequences to score each partial sequence during autoregressive decoding. Meanwhile, Deng & Raffel (2023) used a custom tokenwise loss to distill a reward model trained on full sequences. Thus, it is unclear whether these approaches can give rise to a sound tokenwise text generation policy. Controlled decoding (CD; Mudgal et al., 2024), on the other hand, uses rollouts from the base model along with a reward model trained on full sequences to distill the partial reward.
@@
In this work, we analyze this common RGTG approach. First, we show that the usage of full-sequence reward models to score partial sequences in a tokenwise policy is pathological. To alleviate this, we propose to explicitly train a Bradley-Terry (BT) reward model on partial sequences. We prove that this text generation policy is a ratio of two different RLHF policies trained on sequences of different lengths. Ideally, the policy would be derived from a single RLHF policy, but as we shall also show in Section˜3, the exact computation of such a policy is intractable. By deriving the policy from a ratio of distinct RLHF policies, we obtain a tractable sampling procedure. This is akin to the argument of Zhao et al. (2024) where they use a similar ratio to derive a sequential Monte Carlo method.
@@
    We thus propose to explicitly train a BT reward model on partial sequences and sample from the induced per-token policy induced by it during the decoding time.
@@
    We show that this reward model induces a ratio of two distinct RLHF policies over sequences with different lengths. This is a trade-off that one must make to make tokenwise RGTG free of the aforementioned deficiency and yet still tractable.
@@
    Extensive experiments with multiple LLMs on text generation validate our insights.
@@ 2.1 Reinforcement Learning from Human Feedback
Given a preference dataset $\mathcal{D}=\{({\mathbf{x}}_{k},{\mathbf{y}}_{wk},{\mathbf{y}}_{lk})\}_{k=1}^{K}$ containing $K$ triples of token sequences $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$, Ziegler et al. (2019b) and Ouyang et al. (2022) proposed a technique based on reinforcement learning (RL) to align an LLM with the preference dataset. They train a parametric reward model $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$ that assigns a higher score to the “winning” (i.e., preferred) utterance ${\mathbf{y}}_{w}$ and a lower score to the “losing” utterance ${\mathbf{y}}_{l}$. This is done via the BT model (Bradley & Terry, 1952) which minimizes the loss:
@@
where $\sigma$ is the logistic function. Note that $r_{\phi}$ is trained to score entire utterances $\mathbf{y}$}. Once $r_{\phi}$ is trained, it can be used to infer the probability of generating sequence $\mathbf{y}$} in response to $\mathbf{x}$}, i.e., $P_{\phi}({\mathbf{y}}|{\mathbf{x}})=\nicefrac{{\exp(r_{\phi}({\mathbf{y}}|{\mathbf{x}}))}}{{\sum_{{\mathbf{y}}^{\prime}}\exp(r_{\phi}({\mathbf{y}}^{\prime}|{\mathbf{x}})}}$. Given a reference LLM, we denote by $\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})$ the conditional probability that it will generate response $\mathbf{y}$} to prompt $\mathbf{x}$} (also referred to as policy). We refer to the LLM and its policy interchangeably. One can then copy the LLM and fine-tune it to maximize
@@
where the KL term forms a regularizer that ensures that the fine-tuned model will not differ too much from the reference model. The optimization problem above can be solved by many RL techniques, including the popular PPO algorithm (Schulman et al., 2017). This RL optimization is quite costly in practice due to the size of the LLM.
The optimization (2) has a closed form solution of the form (Peters & Schaal, 2007)
@@
Maximizing the above objective with respect to $\theta$ directly fine-tunes the LLM without the need to learn a reward model. Furthermore, this maximization is done by supervised learning, which is generally simpler than RL. This approach, known as direct preference optimization (DPO, Rafailov et al., 2023), reduced the cost of RLHF while solving an equivalent optimization problem. However, note that both PPO and DPO based RLHF are still very costly in practice since they require fine-tuning (a copy of) the target LLM $\pi_{\text{ref}}$.
@@ 2.2 Reward-Guided Text Generation
However, it is unclear whether the resulting distribution is equivalent/approximating the RLHF policy in (3). Khanov et al. (2024) do train the reward model with the BT loss, but it is trained only with complete sequences, i.e. $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$, while it is used to score partial sequences, i.e. $r_{\phi}({\mathbf{y}}^{1:i}|{\mathbf{x}})$. Hence, it is unclear whether the inferred scores for partial sequences are reasonable. In Section˜3 we show that reward models trained only with complete sequences can assign arbitrary scores to partial sequences, and in Section˜5 we show empirically that the resulting RGTG policy therefore underperforms that of PPO or DPO. Meanwhile, Deng & Raffel (2023) learn the reward model by minimizing a cumulative squared loss to distill a full sequence reward model instead of using the BT loss (1), making the connection to RLHF policy looser. Mudgal et al. (2024) also distill a partial reward model from a full sequence reward model, but the tokenwise policy is not the marginal of the full-sequence policy. Nevertheless, tokens are sampled from a different tokenwise RL formulation that follows a similar derivation as RLHF.
@@
The generated sequences $\mathbf{y}$} are then approximately equal to the sequences generated by the RLHF policy (3). However, their method is general and does not specifically target RGTG—indeed, Zhao et al. (2024) focused on using the implied approximation of the partition function $Z({\mathbf{x}})$. Finally, Rafailov et al. (2024) modifies DPO to obtain a partial-sequence reward model.
@@
Similar to the sequence-based DPO, this reward model is then used to obtain a per-token loss function to fine-tune the LLM and thus, while defining a partial-sequence reward model, is not a RGTG method.
@@ 3 Pitfalls of RGTG and How to Fix Them
First, we start by analyzing the partial sequence rewards inferred from a reward model trained with full sequences only. Proof in Appendix˜A.
@@
A reward model $r$ trained to minimize the BT loss (1) on full sequences ${\mathbf{y}}^{1:|{\mathbf{y}}|}$ may assign arbitrary rewards to partial sequences ${\mathbf{y}}^{1:i}$ (where $i<|{\mathbf{y}}|$). More precisely, $r({\mathbf{y}}^{1:i}|{\mathbf{x}})=v_{{\mathbf{x}},{\mathbf{y}}^{1:i}}$ where $v_{{\mathbf{x}},{\mathbf{y}}^{1:i}}\in\mathbb{R}$ can be any value.
@@
This leads to an unidentifiability problem—see Fig.˜1 for an example. If we learn a reward model based on preferences over full sequences only as proposed by Khanov et al. (2024) and Deng & Raffel (2023), then we may not obtain adequate rewards for partial sequences. As a concrete example, suppose that $r$ is a reward model such that (Fig.˜1)
@@
This reward model satisfies the identity in (8) and therefore, could be the solution when minimizing the BT loss (1). If we use this reward model to sample from the induced RLHF optimal policy in (3), then the token-level sampling distribution is the same as for the reference LLM $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ for all tokens except the last one. This is problematic since RLHF generally changes the token-level distribution at each position, not just the last token. Hence the ARGS method (Khanov et al., 2024) may utilize a reward model that does not score partial sequences properly, and negatively impact token-by-token generation.
Figure 2: The proposed approach to alleviating the problem in Theorem˜1. First, we (i) train the reward model $r$ on partial sequences explicitly (when $|{\mathbf{y}}_{w}|\neq|{\mathbf{y}}_{l}|$, we pad to the longest sequence) and (ii) sample from the weighted sum of the logits and the rewards of the next token during decoding. This is in contrast to some previous RGTG methods where the reward model is trained on full sequences, but decoding relies on partial sequence scoring.
To alleviate this issue, we propose to explicitly train the reward model with partial sequences—still using the BT model in contrast to Deng & Raffel (2023)—as follows (Fig.˜2). We create a separate loss function for all prefix lengths $i$:
@@
Then, given that full sequence ${\mathbf{y}}_{w}$ is preferred to full sequence ${\mathbf{y}}_{l}$, we assume that the partial sequence ${\mathbf{y}}_{w}^{1:i}$ is also preferred to the partial sequence ${\mathbf{y}}_{l}^{1:i}$. Strictly speaking, it is hard for human annotators to compare partial sequences due to their incomplete nature, and most preference datasets do not include preferences over partial sequences. Nevertheless, we can interpret ${\mathbf{y}}_{w}^{1:i}$ as the prefix of a winning sequence that is preferred over a losing sequence with prefix ${\mathbf{y}}_{l}^{1:i}$. The following lemma shows that the resulting reward model ensures that the probability that a first partial sequence is preferred to a second partial sequence corresponds to the probability that the first sequence is extended to a winning full sequence while the second sequence is extended to a losing full sequence according to the preference data distribution $P_{\text{data}}$. Proof in Appendix˜A.
@@
Hence, optimizing the partial-sequence objective (4) for all lengths $i$ determines a reward model for all response prefixes that is adequate in the sense that it induces a distribution over partial sequences that approximates the true underlying preference distribution (due to finite data) instead of assigning arbitrary rewards in the sense of Theorem˜1.
@@
Algorithm 1 summarizes the decoding procedure. Contrary to the previous approach of Khanov et al. (2024), it directly follows the policy induced by the explicitly trained reward model on partial sequences. Meanwhile, compared to Deng & Raffel (2023), it uses the standard BT model instead of a custom squared loss function that distills a full-sequence reward model.
Let us now analyze the tokenwise sampling distribution in (6). By the definition of conditional distributions, we can rewrite it as a ratio of two partial sequence distributions: $\pi(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})=\nicefrac{{\pi({\mathbf{y}}^{1:i}|{\mathbf{x}})}}{{\pi({\mathbf{y}}^{1:i-1}|{\mathbf{x}})}}$. However, it is still unclear how this distribution relates to RLHF policies —- the main point of the tokenwise RGTG methods. The following theorem shows how the decoding process by following this distribution relates to RLHF-induced policies. Proof in Appendix˜A.
@@
Given a reward model trained according to the partial-sequence BT objective in (4), the induced token generation distribution $\pi$ (6) is proportional to the ratio:
@@
where $\pi_{\textnormal{RLHF},i}$ and $\pi_{\textnormal{RLHF},i-1}$ are two distinct policies over prefix sequences of length $i$ and i1i-1, respectively, induced by RLHF optimization (2).
Ideally, we would like a decoding procedure that samples the next token from a distribution that is mathematically equivalent to the conditional distribution resulting from an RLHF over full sequences. However, as shown in Theorem˜3, a partial-sequence reward model $r_{\phi}$ leads to multiple RLHF decoding policies with different conditional distributions for each prefix length $i$. Hence, it is not possible to have equivalence with a single RLHF policy, e.g. as obtained via PPO or DPO.
One may then ask: Which RLHF policy is best? We argue that none of them is necessarily better than the others since they simply arise from considering different prefix lengths. Note that the reward model $r_{\phi}$ leads to a distribution that approximates the true underlying preference distribution on partial sequences. The problem is inherent to RLHF which takes a reference LLM with a consistent distribution over response prefixes induced by a reward model and yields different decoding policies for different prefix lengths.
@@
However, the summations in the above equation are exponentially large in the length $|{\mathbf{y}}|$ of the sequences. This exponential complexity was also noted by Zhao et al. (2024) who proposed a twisted sequential Monte Carlo technique to approximate the computation and mitigate the exponential complexity. In contrast, our approach embraces the multitude of RLHF policies and leverages them in a linear time decoding procedure without any approximation of the partial sequence RLHF policies. The ratio policy (6) described here can thus be seen as a necessary tradeoff if one wants to perform tokenwise RGTG without the pathology in Theorem˜1.